OpenStudy (jacobciezki):

Forty-five is 20% of what number

OpenStudy (jacobciezki):

how you do this

OpenStudy (anonymous):

You can think of the number as x

OpenStudy (anonymous):

so 20%x=45

OpenStudy (jacobciezki):

ok help

OpenStudy (anonymous):

um hold on jake k

OpenStudy (jacobciezki):

k

OpenStudy (anonymous):

20% is the same thing as 0.2, so 0.2x=45. To find x, you divide both sides by 0.2. what's 45/0.2?

OpenStudy (anonymous):

turn the 20% into a decimal, which is .20 now follow the other instructions which make it .20x=45 now all you have to do is divide both sides by .20, and boom x=

OpenStudy (anonymous):

its 225 mate ^_^

OpenStudy (anonymous):

the number 45 is 20% of the number 225

OpenStudy (anonymous):

did get my text jake

OpenStudy (jacobciezki):

yea

OpenStudy (anonymous):

did that help a lot or a little

OpenStudy (jacobciezki):

ok

OpenStudy (anonymous):

if its wrong im srry

OpenStudy (anonymous):

@jacobciezki - just do 20% * 45 because of means multiply.

OpenStudy (jacobciezki):

ok

OpenStudy (anonymous):

Forty-five is 20% of 225. n * 20% = 45; n * 0.2 = 45 n = 45/0.2; n = 225

OpenStudy (jacobciezki):

i got 9

OpenStudy (anonymous):

is this right @jcooper344

OpenStudy (anonymous):

or any one

OpenStudy (jacobciezki):

yea

OpenStudy (anonymous):

ok thx jake

OpenStudy (anonymous):

have to keep my metals

OpenStudy (anonymous):

do still need help jake or no

OpenStudy (jacobciezki):

yes

OpenStudy (anonymous):

hold on ill brb jake ok hold on

OpenStudy (jacobciezki):

ok

OpenStudy (jacobciezki):

wait here is a q

OpenStudy (jacobciezki):

help

OpenStudy (jacobciezki):

no

OpenStudy (jacobciezki):

Sixty-three is 90% of what number?

OpenStudy (jacobciezki):

@*louisa*

OpenStudy (jacobciezki):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? A. B. C. D.

OpenStudy (anonymous):

ok im back jake still need help

OpenStudy (jacobciezki):

yes help

OpenStudy (anonymous):

wts the question

OpenStudy (jacobciezki):

*louisa*

OpenStudy (jacobciezki):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? A. B. C. D.

OpenStudy (anonymous):

is there anwsers

OpenStudy (anonymous):

Ur question isnt coming out right....

OpenStudy (anonymous):

63 is 90% of 70

OpenStudy (jacobciezki):

sryShalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay?

OpenStudy (anonymous):

Not sure on this one sorry :/

OpenStudy (anonymous):

jake she left

OpenStudy (jacobciezki):

wait

OpenStudy (anonymous):

Clearly u just copied and pasted that @WOLFPACK13

OpenStudy (anonymous):

sry louisa im not that great at math ok i said it

OpenStudy (jacobciezki):

9/120=x/100 120/x=100/9 x/120=100/9 120/9=100/x

OpenStudy (jacobciezki):

there

OpenStudy (anonymous):

jake is there any other subject that u need help on besides math

OpenStudy (jacobciezki):

nop just this

OpenStudy (anonymous):

awwwwwwwwwwwwwwwww

OpenStudy (jacobciezki):

what is it

OpenStudy (anonymous):

the reason why i said that is cuz im not good at math other things yes but math is not 1 of them

OpenStudy (jacobciezki):

ok

OpenStudy (anonymous):

im gunna go to english now IM SO SRRY FOR LEAVING YOU NOW

OpenStudy (jacobciezki):

wait